# AsymDex: Leveraging Asymmetry and Relative Motion in Learning Bimanual Dexterity

### v1

@@ Appendix B Grasping Learning
Both AsymDex (our approach) and 2-stage-monolithic policy (one of the baselines in Sec. 4.2) are two-stage policies. Therefore, they can first learn a grasping policy for the facilitating hand (or two grasping policies for facilitating hand and dominant hand respectively). Such policy $\pi_{\mathrm{grasp}}(\hat{\xi}^{h}_{f}(t)|\xi^{h}_{f}(t),\xi^{b}_{f}(t)\circleddash X_{o_{f}})$ takes in the hand joint states and the relative pose between hand and the object, and outputs the target hand joint positions to grasp the object firmly. We first provide pre-grasp annotations [45], which allows the hands to initialize at the position close to the objects with proper joint positions. Then we script the $6D$ lifting hand base motions and design the the following rewards, which is the same across all objects.
@@
Another baseline in Sec. 4.2, i.e., the 1-stage-monolithic policy, does not incorporate the task decomposition. Therefore, it only uses the task-specific interaction rewards (see Appendix. C) to learn how to complete the entire bimanual task. For a fair comparison, both hands also start at the pre-grasp poses.
Appendix C Task Design
In this section, we show the details for each task.
@@ Appendix C Task Design
For each task, the hand joint states $\xi^{h}_{f}(t)$, $\xi^{h}_{d}(t)$ include the 24-DoF hand joint positions and the 24-DoF hand joint velocities. We use quaternions to represent rotation part of object and hand base poses. And for all policies, we also include the previous actions in the policy input. For the block in cup task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the cup and the block respectively. For the stack task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of two cups. For the Bottle cap task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the bottle and the cap respectively. For the Switch task, $o_{f}(t)$ and $o_{d}(t)$ represent the poses of the switch body and the button respectively. The dimensions of the observation and action spaces of each policy are shown in Table. 3. It is obvious that AsymDex policy significantly reduces the state dimensions.
@@
Monolithic Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
@@
  • Place block in cup: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians. The block is initialized in the dominant hand. Thus its position and rotation is calculated based on the initial position and rotation of dominant hand base. The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
@@
  • Stack cups: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians. The cup is initialized in the dominant hand. Thus its position and rotation is calculated based on the initial position and rotation of dominant hand. The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
@@
  • Open bottle cap: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.58,0.62)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.21,-0.19)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.58,0.62)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians. The initial position of the facilitating hand base is randomized: $x_{f}\in\mathcal{X}\sim\mathcal{U}(0.53,0.57)$, $y_{f}\in\mathcal{Y}\sim\mathcal{U}(0.59,0.61)$, $z_{f}\in\mathcal{Z}\sim\mathcal{U}(0.43,0.45)$. For the rotation of the facilitating hand, we randomly rotate it around the axis along the arm at a random angle, $\beta\in\mathcal{B}\sim\mathcal{U}(-0.5,0.5)$, in radians. The bottle is initialized in the facilitating hand. Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
@@
  • Turn on switch: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.25,-0.05)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.5,0.9)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians. The initial position of the facilitating hand base is randomized: $x_{f}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{f}\in\mathcal{Y}\sim\mathcal{U}(0.05,0.25)$, $z_{f}\in\mathcal{Z}\sim\mathcal{U}(0.41,0.81)$. For the rotation of the facilitating hand, we randomly rotate it around the axis along the arm at a random angle, $\beta\in\mathcal{B}\sim\mathcal{U}(-1.0,1.0)$, in radians. The switch is initialized in the facilitating hand. Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
@@
  • Place block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters. This distance makes sure the task is only considered successful when the block is inside the cup. If the block falls on the ground or has not entered the cup within a certain time step, the task is considered failed.
  • Stack cups The task is considered successful if the distance between the cup centers is smaller than 0.02 meters. If either cup falls on the ground or has not been stacked within a certain time step, the task is considered failed.
  • Open bottle cap The task is considered successful if the cap is taken off from its original position 0.05 meters away within a time duration, and is considered failed otherwise.
  • Turn on switch The button and the switch body are connected by a revolute joint ranging from 0 to 0.5585 rads. The task is considered successful if the button is pressed and rotated 0.3585 rads within a time duration, and is considered failed otherwise.
@@
$Reward=R_{hand\ distance}+R_{progress}+R_{action\ penalty}+R_{success\ bonus}$
@@
  • Place block in cup: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth. $R_{progress}=-||x_{cup}-x_{block}||$, where $x_{cup}$ is the position of the cup, and $x_{block}$ is the position of the block.
@@
  • Stack cups: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth, which is grasped by the facilitating hand. $R_{progress}=-||x_{cupd}-x_{cupf}||$, where $x_{cupf}$ is the position of the cup grasped by the facilitating hand, and $x_{cupd}$ is the position of the cup grasped by the dominant hand.
@@
  • Open bottle cap: $R_{hand\ distance}=(1-(||x_{index}-x_{cap}||+||x_{thumb}-x_{cap}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{cap}$ is the position of the bottle cap. $R_{progress}=||x_{cap}-x_{bottle\ top}||$, where $x_{cap}$ is the position of the cap, and $x_{bottle\ top}$ is the position of the top of the bottle.
@@
  • Turn on switch: $R_{hand\ distance}=(1-(||x_{index}-x_{button}||+||x_{thumb}-x_{button}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{button}$ is the position of the button. $R_{progress}=2*\theta_{button}$, where $\theta_{button}$ is the rotated angle of the joint that connects the button and the switch body.

### v2

@@ Appendix B Grasping Learning
Both AsymDex (our approach) and 2-stage-sym policy (one of the baselines in Sec. 5.1.2) are two-stage policies. Therefore, they can first learn a grasping policy for the facilitating hand (or two grasping policies for facilitating hand and dominant hand respectively). Such policy $\pi_{\mathrm{grasp}}(\hat{\xi}^{h}_{f}(t)|\xi^{h}_{f}(t),\xi^{b}_{f}(t)\circleddash P_{f})$ takes in the hand joint states and the relative pose between hand and the object, and outputs the target hand joint positions to grasp the object firmly. We first provide pre-grasp annotations [50], which allows the hands to initialize at the position close to the objects with proper joint positions. Then we script the $6D$ lifting hand base motions and design the the following rewards, which is the same across all objects.
@@
Another baseline in Sec. 5.1.2, i.e., the monolithic policy, does not incorporate the task decomposition. Therefore, it only uses the task-specific interaction rewards (see Appendix. C) to learn how to complete the entire bimanual task. For a fair comparison, both hands also start at the pre-grasp poses.
Appendix C BiDexHands Task Design
In this section, we show the details for each BiDexHands simulation task.
@@ Appendix C BiDexHands Task Design
For each task, the hand joint states $\xi^{h}_{f}(t)$, $\xi^{h}_{d}(t)$ include the 24-DoF hand joint positions and the 24-DoF hand joint velocities. We use quaternions to represent rotation part of object and hand base poses. And for all policies, we also include the previous actions in the policy input. For the block in cup task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the cup and the block respectively. For the stack task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of two cups. For the Bottle cap task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the bottle and the cap respectively. For the Switch task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the switch body and the button respectively. The dimensions of the observation and action spaces of each policy are shown in Table. 4. It is obvious that AsymDex policy significantly reduces the state dimensions.
@@
Sym Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
@@
  • Block in cup: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians. The block is initialized in the dominant hand. Thus its position and rotation is calculated based on the initial position and rotation of dominant hand base. The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
@@
  • Stack: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.3,0.7)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.2,0.0)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.7,1.1)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.57,1.57)$, in radians. The cup is initialized in the dominant hand. Thus its position and rotation is calculated based on the initial position and rotation of dominant hand. The initial position of the facilitating hand base is at $[0.55,0.6,0.8]$.
@@
  • Bottle cap: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.58,0.62)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.21,-0.19)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.58,0.62)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians. The initial position of the facilitating hand base is randomized: $x_{f}\in\mathcal{X}\sim\mathcal{U}(0.53,0.57)$, $y_{f}\in\mathcal{Y}\sim\mathcal{U}(0.59,0.61)$, $z_{f}\in\mathcal{Z}\sim\mathcal{U}(0.43,0.45)$. For the rotation of the facilitating hand, we randomly rotate it around the axis along the arm at a random angle, $\beta\in\mathcal{B}\sim\mathcal{U}(-0.5,0.5)$, in radians. The bottle is initialized in the facilitating hand. Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
@@
  • Switch: The initial position of dominant hand base is randomized: $x_{d}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{d}\in\mathcal{Y}\sim\mathcal{U}(-0.25,-0.05)$, $z_{d}\in\mathcal{Z}\sim\mathcal{U}(0.5,0.9)$. For the rotation of the dominant hand, we randomly rotate it around the axis along the arm at a random angle, $\alpha\in\mathcal{A}\sim\mathcal{U}(-1.0,1.0)$, in radians. The initial position of the facilitating hand base is randomized: $x_{f}\in\mathcal{X}\sim\mathcal{U}(0.2,0.6)$, $y_{f}\in\mathcal{Y}\sim\mathcal{U}(0.05,0.25)$, $z_{f}\in\mathcal{Z}\sim\mathcal{U}(0.41,0.81)$. For the rotation of the facilitating hand, we randomly rotate it around the axis along the arm at a random angle, $\beta\in\mathcal{B}\sim\mathcal{U}(-1.0,1.0)$, in radians. The switch is initialized in the facilitating hand. Thus its position and rotation is calculated based on the initial position and rotation of facilitating hand base.
@@
  • Block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters. This distance makes sure the task is only considered successful when the block is inside the cup. If the block falls on the ground or has not entered the cup within a certain time step, the task is considered failed.
  • Stack The task is considered successful if the distance between the cup centers is smaller than 0.02 meters. If either cup falls on the ground or has not been stacked within a certain time step, the task is considered failed.
  • Bottle cap The task is considered successful if the cap is taken off from its original position 0.05 meters away within a time duration, and is considered failed otherwise.
  • Switch The button and the switch body are connected by a revolute joint ranging from 0 to 0.5585 rads. The task is considered successful if the button is pressed and rotated 0.3585 rads within a time duration, and is considered failed otherwise.
@@
$Reward=\alpha_{1}R_{hand\ distance}+\alpha_{2}R_{progress}+\alpha_{3}R_{action\ penalty}+\alpha_{4}R_{success\ bonus}$
@@
  • Block in cup: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth. $R_{progress}=-||x_{cup}-x_{block}||$, where $x_{cup}$ is the position of the cup, and $x_{block}$ is the position of the block.
@@
  • Stack: $R_{hand\ distance}=e^{-||x_{palm}-x_{cup\ mouth}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ mouth}$ is the position of the cup mouth, which is grasped by the facilitating hand. $R_{progress}=-||x_{cupd}-x_{cupf}||$, where $x_{cupf}$ is the position of the cup grasped by the facilitating hand, and $x_{cupd}$ is the position of the cup grasped by the dominant hand.
@@
  • Bottle cap: $R_{hand\ distance}=(1-(||x_{index}-x_{cap}||+||x_{thumb}-x_{cap}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{cap}$ is the position of the bottle cap. $R_{progress}=||x_{cap}-x_{bottle\ top}||$, where $x_{cap}$ is the position of the cap, and $x_{bottle\ top}$ is the position of the top of the bottle.
@@
  • Switch: $R_{hand\ distance}=(1-(||x_{index}-x_{button}||+||x_{thumb}-x_{button}||))^{3}$, where $x_{index}$ and $x_{thumb}$ are the tip position of index finger and thumb respectively, and $x_{button}$ is the position of the button. $R_{progress}=2*\theta_{button}$, where $\theta_{button}$ is the rotated angle of the joint that connects the button and the switch body.
Appendix D Real-world Task Design
In this section, we present the details for each real-world task.
State Space Design
For each task, the hand joint states $\xi^{h}_{f}(t)$, $\xi^{h}_{d}(t)$ include the 16-DoF hand joint positions. We use quaternions to represent rotation part of object and hand base poses. For the block in cup task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the cup and the block respectively. For the pour task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of two cups. For the twist lid task, $\xi_{f}^{obj}(t)$ and $\xi_{d}^{obj}(t)$ represent the poses of the jar and the lid respectively. The dimensions of the observation and action spaces of each policy are shown in Table. 5. AsymDex policy significantly reduces the state dimensions in real-world tasks consistently.
Sym Asym-w/o-rel Rel-w/o-Asym AsymDex (ours)
Observation $60$ $44$ $53$ $30$
Action $44$ $28$ $38$ $22$
Sampling Procedure
The corresponding simulation environments of real-world tasks include two arms and two hands attached to them. Hence, we randomize the initial position of objects and hand poses by randomizing the initial joint angles of two Kinova arms. Then we initialize the poses of objects according to the initial hand poses.
  • Block in cup: The default initial joint angles of the 7-DoF Kinova arms attached to the dominant hand are $[0.0,0.8,0,\pi/2+0.5,0,-1.3,-\pi/2]$. We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{1,\dots,4\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.2,0.2),\quad\forall i\in\{5,6\}$; $\delta\theta_{7}\in\Theta_{7}\sim\mathcal{U}(-0.3,0.3)$. The default initial joint angles of the 7-DoF Kinova arms attached to the facilitating hand are $[-0.0,0.8,0,\pi/2+0.5,0,-1.3,0]$. We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.0,0.0),\quad\forall i\in\{1,\dots,3\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{4,7\}$.
  • Pour: The default initial joint angles of the 7-DoF Kinova arms attached to the dominant hand are $[0.2,0.8,0,\pi/2+0.5,0,-1.3,-\pi/2]$. We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{1,\dots,4\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.2,0.2),\quad\forall i\in\{5,6\}$; $\delta\theta_{7}\in\Theta_{7}\sim\mathcal{U}(-1.1,0.6)$. The default initial joint angles of the 7-DoF Kinova arms attached to the facilitating hand are $[-0.2,0.8,0,\pi/2+0.5,0,-1.3,0]$. We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.0,0.0),\quad\forall i\in\{1,\dots,3\}$; $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.1,0.1),\quad\forall i\in\{4,7\}$.
  • Twist lid: The default initial joint angles of the 7-DoF Kinova arms attached to the dominant hand are $[0.1,0.5,0,\pi/2+0.55,0.75,-1.5,-\pi/2-0.2]$. We randomized each angle by adding a $\delta\theta_{i}$: $\delta\theta_{i}\in\Theta_{i}\sim\mathcal{U}(-0.0,0.0),\quad\forall i\in\{1,\dots,4\}$; $\delta\theta_{5}\in\Theta_{5}\sim\mathcal{U}(-0.05,0.05)$; $\delta\theta_{6}\in\Theta_{6}\sim\mathcal{U}(-0.2,0.0)$; $\delta\theta_{7}\in\Theta_{7}\sim\mathcal{U}(-0.1,0.1)$. The default initial joint angles of the 7-DoF Kinova arms attached to the facilitating hand are $[-0.1,0.8,0,\pi/2+0.5,-0.7,-1.35,-0.1]$.
Success Criteria
  • Block in cup The task is considered successful if the distance of the block center and the cup center is smaller than 0.035 meters. This distance makes sure the task is only considered successful when the block is inside the cup. If the block falls on the ground or has not entered the cup within a certain time step, the task is considered failed.
  • Pour The task is considered successful if the distance between the rims of the two cups is less than 0.035 meters and the cup grasped by the facilitating hand is up-right. If either cup falls on the ground or the task is not success within a certain time step, the task is considered failed.
  • Twist lid We utilize the articulated bottle simulation of the previous work [19]. The lid and bottle are connected by a single revolute joint. The task is considered successful if the lid is rotated over $3\times\pi$ rad within a time duration, and is considered failed otherwise.
Reward Design
The reward function structure is similar across all tasks:
$Reward=R_{task}+\alpha_{2}R_{action\ penalty}+\alpha_{3}R_{success\ bonus}$
For each task, $R_{action\ penalty}=-||a(t)||^{2}$, and the $R_{success\ bonus}$ is the task success bonus.
  • Block in cup: $R_{task}=\beta_{1}R_{hand\ dist}+\beta_{2}R_{progress}$. $R_{hand\ dist}=e^{-||x_{palm}-x_{cup\ rim}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ rim}$ is the position of the cup rim. $R_{progress}=-||x_{cup}-x_{block}||$, where $x_{cup}$ is the position of the cup, and $x_{block}$ is the position of the block.
  • Pour: $R_{task}=\beta_{1}R_{hand\ dist}+\beta_{2}R_{progress}+\beta_{3}R_{cup\ orient}$. $R_{hand\ dist}=e^{-||x_{palm}-x_{cup\ rim}||}$, where $x_{palm}$ is the dominant hand palm position, and $x_{cup\ rim}$ is the position of the cup rim, which is grasped by the facilitating hand. $R_{progress}=-||x_{cupd}-x_{cupf}||$, where $x_{cupf}$ is the position of the cup rim grasped by the facilitating hand, and $x_{cupd}$ is the position of the cup rim grasped by the dominant hand. $R_{cup\ orient}=\mathbf{z}_{cupf}\cdot\mathbf{z}_{world}$, where $\mathbf{z}_{cupf}$ is the z-axis unit vector of the cup grasped by the facilitating hand, and $\mathbf{z}_{world}$ is the z-axis unit vector of the world frame.
  • Twist lid: $R_{task}=\beta_{1}R_{orient}+\beta_{2}R_{twist}+\beta_{3}R_{finger\ dist}+\beta_{4}R_{hand\ dist\ penalty}$. $R_{orient}=\mathbf{z}_{bottle}\cdot\mathbf{z}_{world}$, where $\mathbf{z}_{bottle}$ is the z-axis unit vector of the bottle grasped by the facilitating hand, and $\mathbf{z}_{world}$ is the z-axis unit vector of the world frame. $R_{twist}=\theta_{lid}^{t}-\theta_{lid}^{t-1}$, where $\theta_{lid}^{t}$ is the current bottle-lid revolute joint angle, and $\theta_{lid}^{t-1}$ is the previous bottle-lid revolute joint angle. And $R_{finger\ dist}$ is the finger contact reward we adopted from a previous lid twisting work [19]. $R_{hand\ dist\ penalty}=-min((||x_{f}-x_{d}||-0.1),0.0)$, where $x_{f}$ and $x_{d}$ are the position of the facilitating hand palm and the dominant hand palm.
Domain Randomization
The domain randomization details are shown in Table. 6
Table 6: Domain Randomization Setup.
Object: Friction (only for twist lid task) [0.5, 1.5]
Hand: Friction (only for twist lid task) [0.5, 1.5]
Object Pos Observation Noise $+\mathcal{N}(0,0.02)$
Hand Joint Observation Noise $+\mathcal{N}(0,0.2)$
Hand Pos Observation Noise $+\mathcal{N}(0,0.02)$
Hand Orientation Observation Noise $+\mathcal{N}(0,0.05)$
Action Noise (only for twist lid task) $+\mathcal{N}(0,0.1)$